OpenStudy (anonymous):

Jacob and Roman are standing next to each other. Jacob, who is one hundred eighty centimeters tall, has a shadow that is thirty centimeters long. If Roman is one hundred sixty-eight centimeters tall, approximately how long is his shadow?

OpenStudy (anonymous):

OK This is a ratio \[\frac{ 180 }{ 30 }\] =\[\frac{ 168 }{ ? }\] So you do cross multiplication 30*168=? then divide ? by 180

OpenStudy (anonymous):

k

OpenStudy (anonymous):

OK So What would your answer be?

OpenStudy (anonymous):

so it would =to 28

OpenStudy (anonymous):

Correcto!

OpenStudy (anonymous):

thanks @JoStudy

OpenStudy (anonymous):

No Problem!